How many unit cells are present in a cube shaped ideal crystal of NaCl of mass of 1g

Grade:12th pass

1 Answers

Vikas TU
3 years ago
Dear student 
Moles of Nacl = 1/58.5 moles 
No. of NaCl molecules = 1/58.5 * 6.023 * 10^23 molecules 
There are 4 molecules of NaCl present in an unit cell
No of unit cell = [6.023/(58.5 *4) ] * 10^23 = 2.57 * 10^21 unit cells 
Good Luck